--- k is in the set of non-negative Real Numbers. PLEASE SHOW ALL WORK. A.) Provide...

T(s) =k \frac{s^{2}-8s+32}{(s^{2})}

--- k is in the set of non-negative Real Numbers.

PLEASE SHOW ALL WORK.

A.) Provide an ACCURATE HAND SKETCHED ROOT LOCUS

B.) Asymptote Intersection:

C.) Asymptote angles :

D.) Breakaway point(s):

E.) Departure/Arrival angles:

F.) Imaginary axis Intercept(s):
Answer #1

The detailed solution of the given question is shown in the image below:

(0) Root Locus sketch 3k OK=oo K=0 K=0 T() = K. 5²85 +32 SP CLOZ = 9285 +32=0 → 5= 8164-4(32) – 4+ C.L.P two at in number at88+32 S +32 m ut K = - dk --TG2-88492) 26 - 5 (25-3)] = 0 as (52.88 +32) 2 => 98-85 +32 = 5(+4) = 5X-88+32 = 82-45 = - 48 =-(F) Imaginary caxis intercept The intercept is obtained from Routh- Hurwitz vriterion, We know, ou Ws) =s&+K(82-85+32) = 0 8P